OpenStudy (anonymous):

if f(x)=x^2-3x and g(x)=√x-1 then f(g(x))

OpenStudy (anonymous):

i ned help

OpenStudy (anonymous):

plug sqrt(x-1) for x in the f(x) equation

OpenStudy (anonymous):

i know to do all that but what i get as an answer is wrong

OpenStudy (anonymous):

what did u get as your answer?

OpenStudy (australopithecus):

f(x)=x^2-3x and g(x)=√x-1 so, ((x-1)^(1/2))^2 - 3(x-1)^(1/2)